A mass of 50g of calcium carbonate is heated, releasing carbon dioxide and calcium oxide. Which of these recorded masses of products is possible, according to the Law ofConservation of Mass?

1
Expert's answer
2021-05-26T07:07:42-0400

According to the reaction:

CaCO3 = CO2 + CaO

From here the mass of carbon dioxide and calcium oxide produced in the reaction are:

m(CO2) = m(CaCO3) × Mr(CO2) / Mr(CaCO3) = 50 g × 44 g/mol / 100 g/mol = 22 g

m(CaO) = m(CaCO3) × Mr(CaO) / Mr(CaCO3) = 50 g × 56 g/mol / 100 g/mol = 28 g


Answer: 22 g and 28 g